1) Given the points (-2,-3) & (4,-4) write an equation for a line which passes through these two points
2) Given the points (t,7) & (-5,3) find the value of t such that the slope of the line containing these points is -8.

Answer by lwsshak3(11628)

1) Given the points (-2,-3) & (4,-4) write an equation for a line which passes through these two points
2) Given the points (t,7) & (-5,3) find the value of t such that the slope of the line containing these points is -8
**
Standard form of an equation for a straight line: y=mx+b, m=slope, b=y-intercept
..
1) slope=∆y/∆x=(y1-y2)/(x1-x2)=rise/run=(-4-(-3))/(4-(-2))=(-4+3)/(4+2)=-1/6
Equation: y=(-1/6)x+b=-x/6+b
solving for b using coordinates of one of the given points(4,-4)
-4=-4/6+b
b=-4+4/6=-24/6+4/6=-20/6=-10/3
Equation of line: y=-x/6-10/32
..
2)slope=∆y/∆x=(3-7)/(-5-t)=-4/(-5-t)=-8
40+8t=-4
8t=-44
t=-44/8=-5.5